# Don’t Retrieve, Navigate: Distilling Enterprise Knowledge into Navigable Agent Skills for QA and RAG

### 16 Apr 2026 v1

@@ Abstract
Retrieval-Augmented Generation (RAG) grounds LLM responses in external evidence but treats the model as a passive consumer of search results: it never sees how the corpus is organized or what it has not yet retrieved, limiting its ability to backtrack or combine scattered evidence. We present Corpus2Skill, which distills a document corpus into a hierarchical skill directory offline and lets an LLM agent navigate it at serve time. The compilation pipeline iteratively clusters documents, generates LLM-written summaries at each level, and materializes the result as a tree of navigable skill files. At serve time, the agent receives a bird’s-eye view of the corpus, drills into topic branches via progressively finer summaries, and retrieves full documents by ID. Because the hierarchy is explicitly visible, the agent can reason about where to look, backtrack from unproductive paths, and combine evidence across branches. On WixQA, an enterprise customer-support benchmark for RAG, Corpus2Skill outperforms dense retrieval, RAPTOR, and agentic RAG baselines across all quality metrics.
@@ 1 Introduction
We evaluate Corpus2Skill on WixQA (Cohen et al., 2025), an enterprise customer-support benchmark designed for evaluating RAG systems over real-world knowledge bases. Corpus2Skill achieves the highest quality across all metrics, outperforming flat retrieval, RAPTOR, and a multi-turn agentic RAG system, while requiring no embedding index or vector database at serve time. These results suggest a complementary design point for knowledge-grounded systems: invest computation at compile time to make the corpus navigable, rather than solely at query time to make it searchable.
@@
    We evaluate on WixQA (Cohen et al., 2025) against five baselines and conduct ablation studies on hierarchy structure, exploration budget, and serving model choice.
@@ 4.2 Main Results
Table 1: Main results on WixQA (200 expert queries). Corpus2Skill achieves the highest quality across all metrics. Quality cells shaded green (darker = better); cost cells shaded red (darker = higher).
@@
BM25 0.342 0.060 0.364 0.119 0.470 0.386 708 0.007
Dense 0.363 0.074 0.382 0.141 0.536 0.450 656 0.008
Hybrid 0.360 0.071 0.380 0.137 0.524 0.410 698 0.008
RAPTOR 0.389 0.109 0.406 0.189 0.675 0.616 1,807 0.012
Agentic 0.388 0.099 0.404 0.186 0.724 0.481 25,807 0.098
Corpus2Skill 0.460 0.137 0.476 0.231 0.729 0.652 53,487 0.172
@@
Table 1 presents the main results. Corpus2Skill achieves the highest scores across all six quality metrics. On Token F1, Corpus2Skill scores 0.460, a 19% relative improvement over Agentic (0.388) and 27% over Dense retrieval (0.363). On Factuality, Corpus2Skill (0.729) exceeds both Agentic (0.724) and RAPTOR (0.675). On Context Recall, Corpus2Skill leads by a clear margin (0.652 vs. 0.616 for RAPTOR and 0.481 for Agentic), indicating that the navigation-based approach surfaces more relevant content than either vector search or multi-turn agentic retrieval.
@@
Both RAPTOR and Corpus2Skill substantially outperform the three flat retrieval baselines (BM25, Dense, Hybrid), confirming that hierarchical document organization improves answer quality. RAPTOR’s collapsed tree retrieval achieves Factuality of 0.675 and Context Recall of 0.616, substantially above flat methods (Dense: 0.536/0.450, BM25: 0.470/0.386). Cluster summaries surface thematic connections that flat embeddings miss, allowing the retrieval system to match queries to relevant document groups rather than individual passages.
@@ 4.4 Ablation Studies
We conduct three ablation studies to characterize how compile-time and serve-time design choices affect the quality-cost tradeoff.
@@
Table 2: Ablation: cluster structure. Narrow trees match or exceed the default; wide trees degrade quality.
@@
Narrow (p=5p{=}5) 0.461 0.736 0.674 0.186
Default (p=10p{=}10) 0.460 0.729 0.652 0.172
Wide (p=20p{=}20) 0.361 0.410 0.355 0.242
We vary the branching ratio $p$ and top-cluster count $K$ to produce hierarchies of different shapes (Table 2). With p=5p{=}5, the narrow branching ratio creates a 4-level hierarchy with 3 top-level skills. With deeper hierarchies or larger corpora, the file count per skill can exceed the API’s 200-file limit. We introduce a compact mode that merges near-leaf INDEX.md files into their parent: instead of creating a separate subdirectory for each lowest-level group, the parent’s INDEX.md contains labeled sections listing each sub-group’s documents inline. This reduces file count by up to 80% while preserving all navigational information. The narrow hierarchy slightly outperforms the default on all quality metrics (F1 0.461 vs. 0.460, Factuality 0.736 vs. 0.729, Context Recall 0.674 vs. 0.652) at a modest 8% cost increase (\$0.186 vs. \$0.172 per query). The deeper hierarchy provides more fine-grained topic groupings, and the compact merged files give the agent a denser view of the leaf-level documents in fewer navigation steps. In the wide-shallow variant (p=20p{=}20), F1 drops 21% and Factuality drops 44% while costing 41% more. With only 2 extremely broad skills covering approximately 3,000 documents each, the SKILL.md summaries become too generic for effective topic routing.
@@
Table 3: Ablation: agent exploration budget. Quality is stable across budgets, indicating effective hierarchy organization.
@@
5 0.453 0.721 0.636 0.170
10 0.461 0.748 0.667 0.172
20 0.460 0.729 0.652 0.172
We vary the maximum allowed interaction rounds while keeping the default skill tree (Table 3). The exploration budget has minimal impact on quality: even a budget of 5 achieves F1 of 0.453, only 1.5% below the full-budget configuration. Factuality is nearly identical across variants (0.721 to 0.748), and the marginal cost difference is negligible. The hierarchical structure is sufficiently well-organized that the agent rarely needs extended exploration.
@@
Table 4: Ablation: serving LLM choice. A cheaper model achieves comparable quality at roughly half the cost.
@@
Claude Sonnet 4.6 0.460 0.729 0.652 0.172
Claude Haiku 4.5 0.423 0.645 0.705 0.088
We swap only the serving model while keeping the same compiled skill tree (Table 4). A smaller, cheaper model achieves comparable quality at roughly half the cost (\$0.088 vs. \$0.172 per query). While F1 drops by 8% and Factuality by 12%, Context Recall increases from 0.652 to 0.705, suggesting the cheaper model navigates the hierarchy more thoroughly, retrieving more relevant documents even if its answer synthesis is slightly less polished. At \$0.088 per query, the cheaper model offers a compelling cost-quality tradeoff. This also demonstrates that the compiled skill tree is robust to the choice of serving model: the quality of the hierarchy, not the sophistication of the navigator, is the primary driver of retrieval performance.

### 29 Apr 2026 v2

@@ Abstract
Retrieval-Augmented Generation (RAG) grounds LLM responses in external evidence but treats the model as a passive consumer of search results: it never sees how the corpus is organized or what it has not yet retrieved, limiting its ability to backtrack or combine scattered evidence. We present Corpus2Skill, which distills a document corpus into a hierarchical skill directory offline and lets an LLM agent navigate it at serve time. The compilation pipeline iteratively clusters documents, generates LLM-written summaries at each level, and materializes the result as a tree of navigable skill files. At serve time, the agent receives a bird’s-eye view of the corpus, drills into topic branches via progressively finer summaries, and retrieves full documents by ID. Because the hierarchy is explicitly visible, the agent can reason about where to look, backtrack from unproductive paths, and combine evidence across branches. On WixQA, an enterprise customer-support benchmark for RAG, Corpus2Skill outperforms dense retrieval, RAPTOR, and agentic RAG baselines across all quality metrics. We further evaluate generalization on nine RAGBench subsets reformulated as retrieval-stress benchmarks: Corpus2Skill attains the highest macro-average F1 across the full 10-dataset suite and characterizes a clear regime — single-domain, atomic-document corpora — where corpus navigation is the right primitive, while flat retrieval remains preferable for open-domain or extractive pools.
@@
Code: https://github.com/dukesun99/Corpus2Skill
@@ 1 Introduction
We evaluate Corpus2Skill on WixQA (Cohen et al., 2025), an enterprise customer-support benchmark designed for evaluating RAG systems over real-world knowledge bases. Corpus2Skill achieves the highest quality across all metrics, outperforming flat retrieval, RAPTOR, and a multi-turn agentic RAG system, while requiring no embedding index or vector database at serve time. We complement this anchor result with a generalization study on nine RAGBench (Friel et al., 2024) subsets reformulated as retrieval benchmarks: Corpus2Skill attains the highest macro-average F1 over the full ten-dataset benchmark, and the per-dataset dispersion isolates the regime where navigation is the right primitive (single-domain, atomic-document corpora) versus the regime where flat retrieval remains preferable (open-domain abstracts, very long extractive documents). These results suggest a complementary design point for knowledge-grounded systems: invest computation at compile time to make the corpus navigable, rather than solely at query time to make it searchable.
@@
    We evaluate on WixQA (Cohen et al., 2025) against five baselines, conduct ablation studies on hierarchy structure, exploration budget, and serving model choice, and run a 9-subset generalization study on RAGBench (Friel et al., 2024) that characterizes which corpus shapes Corpus2Skill helps and which it does not.
@@ 4.2 Main Results
[…]
@@
BM25 0.342 0.060 0.364 0.119 0.470 0.386 1.00 708 0.007
Dense 0.363 0.074 0.382 0.141 0.536 0.450 1.00 656 0.008
Hybrid 0.360 0.071 0.380 0.137 0.524 0.410 1.00 698 0.008
RAPTOR 0.389 0.109 0.406 0.189 0.675 0.616 1.00 1,807 0.012
Agentic 0.381 0.098 0.395 0.183 0.719 0.498 5.21 15,483 0.088
Corpus2Skill 0.468 0.143 0.483 0.238 0.739 0.673 2.43 12,601 0.089
@@
Table 1 presents the main results. Corpus2Skill achieves the highest scores across all six quality metrics. On Token F1, Corpus2Skill scores 0.468, a 23% relative improvement over Agentic (0.381) and 29% over Dense retrieval (0.363). On Factuality, Corpus2Skill (0.739) exceeds both Agentic (0.719) and RAPTOR (0.675). On Context Recall, Corpus2Skill leads by a clear margin (0.673 vs. 0.616 for RAPTOR and 0.498 for Agentic), indicating that the navigation-based approach surfaces more relevant content than either vector search or multi-turn agentic retrieval.
@@
Both RAPTOR and Corpus2Skill substantially outperform the three flat retrieval baselines (BM25, Dense, Hybrid), confirming that hierarchical document organization improves answer quality. RAPTOR’s collapsed tree retrieval achieves Factuality of 0.675 and Context Recall of 0.616, substantially above flat methods (Dense: 0.536/0.450, BM25: 0.470/0.386), and Corpus2Skill exceeds RAPTOR by a further 10% on Factuality (0.739) and 9% on Context Recall (0.673). Cluster summaries surface thematic connections that flat embeddings miss, allowing the retrieval system to match queries to relevant document groups rather than individual passages.
@@ 4.3 Case Study: Navigation in Action
4.4 Generalization to RAGBench
Table 2: Generalization study on 9 RAGBench subsets, 200 queries each (or all available when smaller), evaluated on the expanded retrieval pool (union of train+val+test documents, deduplicated by content hash). All methods use Claude Sonnet 4.6 with Anthropic ephemeral prompt caching. Quality cells shaded green (darker = better) and \$/q cells shaded red (darker = higher); the per-dataset best value of each quality metric is shown in bold. Datasets are grouped by the structural pattern that determines whether navigation helps; see Table 3 for the full 10-dataset macro-average. “Inp.” is per-query input tokens billed at the full input rate (in thousands); “Turns” counts outer LLM round-trip API calls only (server-side tool invocations within a single request, e.g. Skills / code-execution skill-file reads, are not counted).
Dataset Method Quality Metrics Cost
[…]
(a) Atomic, single-domain corpora — Corpus2Skill wins
Emanual BM25 0.413 0.108 0.435 0.195 0.658 0.552 1.00 0.54 0.006
Dense 0.423 0.116 0.444 0.205 0.705 0.656 1.00 0.50 0.006
Hybrid 0.429 0.118 0.451 0.209 0.727 0.659 1.00 0.57 0.006
Agentic 0.456 0.147 0.475 0.245 0.836 0.624 3.32 9.0 0.044
Corpus2Skill 0.497 0.160 0.514 0.279 0.794 0.762 3.11 7.2 0.054
DelucionQA BM25 0.413 0.149 0.431 0.231 0.727 0.688 1.00 0.68 0.006
Dense 0.438 0.167 0.458 0.257 0.812 0.740 1.00 0.65 0.006
Hybrid 0.432 0.161 0.451 0.250 0.774 0.708 1.00 0.67 0.006
Agentic 0.441 0.178 0.459 0.272 0.838 0.723 2.83 6.4 0.037
Corpus2Skill 0.462 0.174 0.481 0.285 0.782 0.714 2.71 8.8 0.062
HAGRID BM25 0.325 0.109 0.334 0.195 0.831 0.831 1.00 0.66 0.005
Dense 0.317 0.104 0.326 0.190 0.839 0.881 1.00 0.66 0.005
Hybrid 0.328 0.112 0.337 0.198 0.834 0.858 1.00 0.67 0.005
Agentic 0.365 0.133 0.374 0.223 0.841 0.842 2.31 4.2 0.022
Corpus2Skill 0.417 0.165 0.427 0.262 0.824 0.759 2.17 8.5 0.071
(b) Heterogeneous or table-heavy corpora — flat retrieval competitive, Corpus2Skill close
CovidQA BM25 0.203 0.048 0.212 0.100 0.547 0.548 1.00 0.67 0.006
Dense 0.214 0.049 0.224 0.107 0.643 0.645 1.00 0.64 0.006
Hybrid 0.218 0.053 0.227 0.111 0.623 0.586 1.00 0.66 0.006
Agentic 0.200 0.052 0.206 0.107 0.635 0.514 4.05 10 0.056
Corpus2Skill 0.206 0.044 0.217 0.088 0.452 0.396 9.39 17 0.174
ExpertQA BM25 0.310 0.044 0.337 0.086 0.490 0.433 1.00 0.70 0.007
Dense 0.330 0.050 0.358 0.095 0.580 0.526 1.00 0.71 0.007
Hybrid 0.324 0.049 0.352 0.093 0.542 0.487 1.00 0.71 0.007
Agentic 0.289 0.042 0.310 0.092 0.685 0.435 4.44 14 0.085
Corpus2Skill 0.288 0.039 0.311 0.077 0.448 0.295 2.83 9.0 0.088
TechQA BM25 0.381 0.082 0.406 0.150 0.527 0.310 1.00 0.95 0.009
Dense 0.380 0.087 0.404 0.154 0.565 0.303 1.00 0.97 0.009
Hybrid 0.383 0.083 0.409 0.152 0.567 0.304 1.00 0.96 0.009
Agentic 0.297 0.069 0.313 0.132 0.606 0.291 6.36 28 0.151
Corpus2Skill 0.328 0.067 0.350 0.123 0.424 0.281 4.34 27 0.293
FinQA BM25 0.354 0.108 0.364 0.177 0.341 0.362 1.00 0.73 0.005
Dense 0.401 0.122 0.414 0.206 0.525 0.517 1.00 0.74 0.006
Hybrid 0.390 0.118 0.401 0.198 0.446 0.434 1.00 0.73 0.006
Agentic 0.367 0.114 0.375 0.199 0.566 0.493 4.44 15 0.078
Corpus2Skill 0.382 0.132 0.389 0.217 0.555 0.495 6.25 50 0.491
(c) Homogeneous-tabular or long-extractive corpora — Corpus2Skill loses
TatQA BM25 0.307 0.098 0.317 0.166 0.464 0.468 1.00 0.55 0.005
Dense 0.304 0.086 0.313 0.154 0.519 0.512 1.00 0.63 0.006
Hybrid 0.315 0.094 0.325 0.165 0.537 0.518 1.00 0.59 0.005
Agentic 0.351 0.127 0.359 0.213 0.614 0.546 3.04 8.7 0.045
Corpus2Skill 0.246 0.076 0.253 0.135 0.402 0.363 8.43 80 0.969
CUAD BM25 0.314 0.034 0.338 0.082 0.280 0.191 1.00 0.72 0.007
Dense 0.302 0.033 0.325 0.083 0.282 0.170 1.00 0.70 0.008
Hybrid 0.308 0.034 0.331 0.085 0.277 0.204 1.00 0.71 0.007
Agentic 0.249 0.035 0.265 0.084 0.260 0.193 4.29 13 0.079
Corpus2Skill 0.211 0.028 0.227 0.060 0.229 0.201 8.82 55 0.369
Table 3: Macro-average over the full 10-dataset benchmark (WixQA anchor + 9 RAGBench subsets, Table 2). Per-metric best is in bold. Corpus2Skill leads on every lexical-overlap metric; Agentic wins Factuality (it falls back to “I don’t know” more often, which the judge rewards), and Dense wins Context Recall (which rewards span overlap that embedding retrieval directly optimizes for).
[…]
BM25 0.336 0.084 0.354 0.150 0.534 0.477 1.00 0.006
Dense 0.347 0.089 0.365 0.159 0.601 0.540 1.00 0.007
Hybrid 0.349 0.089 0.367 0.160 0.585 0.517 1.00 0.007
Agentic 0.339 0.099 0.353 0.175 0.660 0.516 4.03 0.068
Corpus2Skill 0.350 0.103 0.378 0.177 0.565 0.494 5.04 0.266
To probe how Corpus2Skill behaves outside its target domain, we run the same five methods on nine subsets of RAGBench (Friel et al., 2024): CovidQA (Möller et al., 2020), DelucionQA (Sadat et al., 2023), Emanual (Nandy et al., 2021), ExpertQA (Malaviya et al., 2024), TechQA (Castelli et al., 2020), FinQA (Chen et al., 2021), CUAD (Hendrycks et al., 2021), HAGRID (Kamalloo et al., 2023), and TatQA (Zhu et al., 2021), sampling 200 test queries per subset (or all available when smaller). RAGBench is published as a generator-evaluation benchmark with per-row gold contexts; to convert it into a retrieval benchmark we build the expanded retrieval pool as the union of all train+val+test documents, deduplicated by content hash. This is intentionally harder than the original RAGBench protocol: every test query must compete against thousands of distractors, and there is no curated retrieval corpus to fall back on. Pool sizes range from 221 (Emanual) to 12,727 (TatQA).
Headline result.
Corpus2Skill attains the highest macro-average F1 (0.350) across the 10-dataset benchmark (Table 3), narrowly ahead of Hybrid (0.349) and Dense (0.347), and it leads on every lexical-overlap metric (BLEU, ROUGE-1, ROUGE-2). Agentic continues to win Factuality (0.660), and Dense wins Context Recall (0.540) because the LLM-judged CtxR rewards span overlap that embedding retrieval is explicitly optimized for. The picture is therefore not a uniform sweep — Corpus2Skill wins on 4 of 10 datasets by a margin of 0.02–0.10 F1, ties on 2, and loses on 4 (Table 2) — but the per-dataset dispersion lets us characterize which corpora navigation actually helps.
When does corpus navigation help?
Three structural properties consistently separate wins from losses:
(i) Atomic, topically-coherent documents. On WixQA, Emanual, DelucionQA, and HAGRID — corpora where each document covers a single feature, entity, or how-to — Corpus2Skill beats every retrieval baseline by 0.02–0.09 F1. HAGRID is particularly informative: even though its corpus is open-domain Wikipedia, every article is by construction a coherent unit about one entity, and the cluster summarizer recovers a clean per-topic taxonomy. The hierarchy gives the agent useful disambiguation that flat retrievers cannot exploit.
(ii) Document granularity in the 0.5–3k character range. The wins cluster around documents long enough to contain a complete answer ($\geq 500$ characters) but short enough that an LLM-written summary captures the article ($\leq 3000$ characters). On CUAD (legal contracts averaging $\approx$7,500 characters of mixed clauses) and TechQA (IBM technotes averaging 3,740 characters), single documents straddle multiple clause types or symptom categories, so any single skill placement is provably suboptimal and Corpus2Skill loses to flat retrieval.
(iii) A recoverable implicit taxonomy. On TatQA, which is essentially homogeneous SEC-style financial-statement tables, the level-1 cluster summaries collapse onto near-identical labels (“corporate financial reporting”) and the agent can no longer distinguish branches at the top level. Corpus2Skill hits max_turns on a third of queries and the cost balloons to \$0.97/query, the worst case in our evaluation. ExpertQA exhibits a milder version of the same failure mode: queries span dozens of unrelated scientific fields, summaries become generic, and the navigator cannot exploit them. Agentic search wins both subsets, presumably because the lack of a useful hierarchy makes the search-then-read pattern strictly more efficient.
Where flat retrieval is the right default.
On CovidQA, the corpus is short, near-duplicate research abstracts whose information density is already saturated by top-5 retrieval; on TechQA, lexical / dense baselines are within a few percent of each other and Corpus2Skill pays a $30\times$ cost premium for no quality gain. The pattern is consistent: when the bottleneck for a corpus is retrieval recall rather than topic disambiguation, hierarchical navigation adds turns and tokens without helping. We therefore position Corpus2Skill as a complementary primitive — strongly preferred for curated, single-domain enterprise knowledge bases (the deployment target this paper is motivated by), with flat retrieval or RAPTOR retained as the right default for open-domain or homogeneous pools.
Competitive cases.
CovidQA and FinQA sit between the wins and the failures. On CovidQA the lexical metrics agree to within 0.02 F1 across all five methods. On FinQA, Corpus2Skill matches the best baseline on lexical n-gram quality (BLEU 0.132 and R-2 0.217 are both highest, F1 within 0.02 of Dense), confirming that financial reports still admit a useful by-firm and by-statement-section hierarchy; the cost premium ($\sim 80\times$ over Dense) rather than the quality is what makes it a poor fit. We treat FinQA and CovidQA as ties rather than losses in our headline count.
Failure modes exposed by the generalization sweep.
The CUAD and TatQA results are the two cleanest cases in our evaluation where the design hypothesis behind Corpus2Skill — that an LLM-written hierarchy provides useful navigation signal — does not hold. CUAD answers are long verbatim quotes from a single contract clause, which token-F1 penalizes for any paraphrase; TatQA is structurally so flat that the cluster summarizer has nothing discriminative to surface. Both motivate concrete follow-on work — span-level extraction tools at the leaf level for the former, table-aware schema-based clustering for the latter — discussed alongside other limitations in Section 5.
@@ 4.5 Ablation Studies
We conduct three ablation studies on the WixQA anchor to characterize how compile-time and serve-time design choices affect the quality-cost tradeoff.
@@
Table 4: Ablation: cluster structure (200 queries, prompt caching on). All three shapes land within $\approx$2 F1 / 2 Fact points; the default shape is cheapest.
@@
Narrow (p=5p{=}5, 3 skills, compact) 0.457 0.734 0.649 0.101
Default (p=10p{=}10, 6 skills) 0.468 0.739 0.673 0.089
Wide (p=20p{=}20, 2 skills) 0.469 0.749 0.666 0.096
We vary the branching ratio $p$ and top-cluster count $K$ to produce hierarchies of different shapes (Table 4). With p=5p{=}5, the narrow branching ratio creates a 4-level hierarchy with 3 top-level skills; with p=20p{=}20 we obtain a wide-shallow tree with only 2 very broad top-level skills that each span roughly 3,000 documents. With deeper hierarchies or larger corpora, the file count per skill can exceed the API’s 200-file limit. We introduce a compact mode that merges near-leaf INDEX.md files into their parent: instead of creating a separate subdirectory for each lowest-level group, the parent’s INDEX.md contains labeled sections listing each sub-group’s documents inline. This reduces file count by up to 80% while preserving all navigational information. All three shapes land within $\approx$2 F1 points and $\approx$2 Factuality points of each other (F1 0.457 / 0.468 / 0.469, Factuality 0.734 / 0.739 / 0.749); the default shape is cheapest at \$0.089 per query, with narrow 14% more expensive (\$0.101) and wide 8% more expensive (\$0.096). The takeaway is that once the per-level LLM-written summaries are faithful, the agent is robust to the tree’s branching factor: wider trees with fewer, broader top-level skills still navigate effectively because INDEX.md files at deeper levels restore the needed specificity, and narrower trees reduce the chance of top-level routing errors but add one extra navigation step per query. These results corroborate our main finding that the quality of the summaries at each level, not the precise shape of the tree, is the primary driver of navigation quality.
@@
Table 5: Ablation: agent exploration budget (200 queries, prompt caching on). Quality is stable across budgets, indicating effective hierarchy organization.
@@
5 0.458 0.741 0.657 0.087
10 0.464 0.737 0.670 0.095
20 0.468 0.739 0.673 0.089
We vary the maximum allowed interaction rounds while keeping the default skill tree (Table 5). The exploration budget has minimal impact on quality: even a budget of 5 achieves F1 of 0.458, just 2% below the full-budget configuration, and Factuality is nearly identical across variants (0.737–0.741). The hierarchical structure is sufficiently well-organized that the agent rarely needs extended exploration. Budget-5 also happens to be the cheapest setting (\$0.087 vs. \$0.089 for budget-20), because shorter traces generate and cache fewer intermediate turns; this makes max_turns=5 an attractive default for cost-sensitive deployments.
@@
Table 6: Ablation: serving LLM choice (200 queries, prompt caching on). Haiku retains 91% of Sonnet’s F1 and matches its Context Recall at 71% of the cost.
@@
Claude Sonnet 4.6 0.468 0.739 0.673 0.089
Claude Haiku 4.5 0.424 0.662 0.698 0.063
We swap only the serving model while keeping the same compiled skill tree (Table 6). Swapping Sonnet 4.6 for the smaller and cheaper Haiku 4.5 reduces per-query cost from \$0.089 to \$0.063 (a 29% reduction) while retaining 91% of Sonnet’s F1 (0.424 vs. 0.468) and 90% of its Factuality (0.662 vs. 0.739). Context Recall actually increases from 0.673 to 0.698, suggesting that Haiku navigates the hierarchy more thoroughly—retrieving more relevant documents—even if its answer synthesis is slightly less polished. At \$0.063 per query, Haiku offers a compelling cost-quality tradeoff for budget-constrained deployments. This also demonstrates that the compiled skill tree is robust to the choice of serving model: the quality of the hierarchy, not the sophistication of the navigator, is the primary driver of retrieval performance.